# Mean-field dynamics of attractive resource interaction: From uniform to aggregated states

### v1

@@ 2. $2)$
    The set of vectors ${\bf p}^{0}$, where at least one coordinate is zero is an invariant set of the system. All fixed points $\bf p$$^{\infty}$ that belong to this set are unstable repellers.
@@
The stability analysis of fixed points for nonlinear maps, as performed here via linearization and eigenvalue analysis of the Jacobian, is a cornerstone of dynamical systems theory [21, 22]. The result that boundary equilibria are unstable repellers while the interior equilibrium is a stable attractor is a common feature in systems modeling population dynamics and competition [23].
@@
In addition to establishing stability, the analysis of the Jacobian’s eigenvalues provides an important insight into the rate of convergence to the equilibrium state. As shown, all n1n-1 eigenvalues corresponding to perturbations on the simplex are equal to $\lambda=\frac{n^{2}-2}{n^{2}-1}$. This value can be expressed as $\lambda=1-\frac{1}{n^{2}-1}$, from which it is evident that as the number of system components $n$ increases, $\lambda$ approaches 1 from below. Since the decay rate of perturbations near the fixed point is determined by the magnitude $|\lambda|^{t}$, the proximity of $\lambda$ to unity signifies that the convergence is very slow. Consequently, for systems with a large number of interacting components (a large $n$), the process of reaching a uniform distribution can be extremely prolonged. In practice, this means the system can spend long periods in transient, non-uniform states before reaching its asymptotic equilibrium.
3 Dependence of the limit state on the set of constants.
@@
This heterogeneity in the parameters $c_{i}$ is conceptually similar to the introduction of influential agents or ”zealots” in opinion dynamics, who can prevent a full consensus [24]. Whereas models of bounded confidence achieve fragmentation through interaction thresholds [20], our model generates aggregated states analytically, based on the intrinsic ’attractiveness’ $c_{i}$ of each component.
@@
Next, let’s consider the case where all parameters $c_{i}$ are different. They specifically influence the resource distribution and lead to an increase or decrease in a component’s share in certain states $\omega_{i}$. The analysis of how such external parameters (assistance or favorability) determine the characteristics of the final equilibrium state is a relevant area of study in these models [25]. Indeed, the limit state is entirely determined by the set of parameters $c_{i}$. For instance, a violation of condition (11) for some $i$ in Theorem 3 leads to $p_{i}^{\infty}=0$, which can be interpreted as the complete depletion of the resource from this $i$-th state. This has been confirmed by numerous computer simulations (see Fig. 2). And the explicit form of the limit vectors demonstrates that the greater the value of the constant $c_{i}$, the greater the value of the corresponding $i$-th limit coordinate. That is, more favorable conditions associated with a certain state lead to an increase in the corresponding component’s share in that state.
@@
Hereafter, we will use the notation $\kappa_{i}^{t}:=c^{t}_{i}(1-p_{i}^{t})$.
@@
$V_{n}=\{{\bf e}_{1},{\bf e}_{2},\dots,{\bf e}_{n}\},$
where ${\bf e}_{i}=(0,\dots,1,\dots,0)$ is the vector with a 1 in the $i$-th position and zeros elsewhere.
@@
Let the initial state of the system be ${\bf p}^{0}={\bf e}_{i}$ for some $i\in\{1,\dots,n\}$. This means the initial components of the system are $p_{i}^{0}=1$ and $p_{j}^{0}=0$ for all $j\neq i$.
@@
3.1 Existence of limit fixed point
@@
Let ${\bf p}\in\mathbb{R}_{+}^{2}$, ${\bf p}^{0}\neq{\bf e}_{i}$. Then each trajectory of the dynamic conflict system (1) converges to a stable limit state, which is a fixed point
@@
Then from the invariance of the limit state, we have the system
@@
Next, the case of a three-dimensional space is considered, i.e., ${\bf p}\in\mathbb{R}_{+}^{3}$. Note that Theorem 2, as well as the following Theorem 3, can be generalized to the case ${\bf p}\in\mathbb{R}_{+}^{n}$, $n>3$. In this case, condition (11) becomes significantly more complex, and the description of the limit vector $\bf p$$^{\infty}$ breaks down into several non-trivial cases that require additional analysis.
a)
b)
Figure 2: Trajectories of the dynamic conflict system for the three-dimensional space case under the action of the conflict transformation (9). Dynamics of change of arbitrary stochastic vectors ${\bf{p}}=(p_{1};p_{2};p_{3})$ with a set of constants: a) ${\bf{c}}=(0.3;0.4;0.25)$, ${\bf{p}}^{\infty}=(0.322;0.4915;0.1865)$ and b) ${\bf{c}}=(0.8;0.1;0.9)$, ${\bf{p}}^{\infty}=(0.4706;0;0.5294)$}. In case b), the condition is violated, i.e., $c_{2}<\frac{c_{1}c_{3}}{c_{1}+c_{3}}$}, and therefore $p_{2}^{\infty}=0$.
@@
Let ${\bf{p}}\in\mathbb{R}_{+}^{3}$, ${\bf p}^{0}\neq{\bf e}_{i}$, and for each $i=\overline{1,3}$ one of the conditions $\kappa^{t}_{i}>L_{c}^{t}$ or $\kappa^{t}_{i}<L_{c}^{t}$ holds for any $t$. Then every trajectory of the dynamic conflict system (1) converges to a limit state, which is a fixed point
@@
This follows directly from equations (9). If this condition holds for all $t$, then the sequence $\{p_{i}^{t}\}_{t=0}^{\infty}$ is monotonic. The boundedness of this sequence follows from the problem statement. Therefore, a limit vector $\bf{p}$$^{\infty}$ exists. Assume all limit coordinates $p_{i}^{\infty}$ are non-zero. Then the invariance of the limit state means that $k_{i,c}^{t}=1$ for all $i=1,2,3$. That is, we have the system
@@
It should be noted that the limit state is stable, as it is determined only by a fixed set of constant parameters $c_{i}$, which in turn are given at $t=0$ and do not depend on time $t$.
@@
Let $M^{*}=\{k:p_{k}^{\infty}>0\}$ be the set of indices for which the coordinates of the limiting vector are strictly positive. Then $\gamma(M^{*})$ is the number of elements in the set $M^{*}$. The existence and uniqueness of $M^{*}$ are determined by the iterative procedure (1). Let’s denote
@@
Let ${\bf{p}}\in\mathbb{R}^{n}$, $n>3$. Suppose that for each $i=\overline{1,n}$, starting from some time $t$, one of the conditions $k_{i}^{t}>L_{c}^{t}$ or $k_{i}^{t}<L_{c}^{t}$ holds. Then every trajectory of the dynamic conflict system (1) converges to a limit state, which is a fixed point
@@
For any $i=1,\dots,n$, one of the following equalities holds:
@@
Similarly to the proof of Theorem 3 for $n=3$, based on the monotonicity of the sequences $\{p_{i}^{t}\}$, the trajectories ${\bf p}^{t}$ converge to a limiting state $p^{\infty}$, which is a fixed point of the system.
For a fixed point $p^{\infty}$, if $p_{i}^{\infty}>0$ (i.e., $i\in M^{*}$) and $c_{i}>0$, it follows from the dynamical equation (9) $p_{i}^{t+1}=p_{i}^{t}\cdot k_{i,c}^{t}$, where $k_{i,c}^{t}=\frac{n-1+c_{i}(1-p_{i}^{t})}{n-1+L_{c}^{t}}$, that in the stationary state $k_{i,c}^{\infty}=1$. This means that the numerator and denominator of $k_{i,c}^{\infty}$ are equal:
@@
The sum of all coordinates of the vector $p^{\infty}$ is equal to 1. Since $p_{k}^{\infty}=0$ for $k\notin M^{*}$, then
@@
For $j\notin M^{*}$, $p_{j}^{\infty}=0$, which holds under the condition $c_{j}\leq\Lambda(M^{*})$. Therefore, the explicit form of the limiting state depends on the set $M^{*}$, which, in turn, is determined by the entire set of constants $c_{k}$. ∎
@@
holds for all $k=\overline{1,n}$, then $\gamma(M^{*})=n$,
@@
As the mean-field vector ${\bf r}^{t}$ is not an independent variable, being defined at each step as a function of ${\bf p}^{t}$ by formula (3), its limit state $\bf r$$^{\infty}$ is also entirely dependent on $\bf p$$^{\infty}$. Therefore, after the limit coordinates $p_{i}^{\infty}$ have been determined from Theorems 24, the corresponding coordinates $r_{i}^{\infty}$ can be readily computed as:
@@
Below, we present the algorithm that allows for the analytical computation of the limit fixed point $\bf p$$^{\infty}$ for a given initial vector ${\bf p}^{t=0}$ and set of control parameters $c_{i}$, $i=\overline{1,n}$. The procedure consists of three main steps: determining the initial set of active coordinates, iteratively finding the final set of non-zero coordinates $M^{*}$, and finally calculating the values of these coordinates. The complete implementation of this algorithm, available as find_fixed_point.py, can be found in the public repository (see Code Availability)
1. Determine the initial set of active indices ($M_{0}$). The dynamics of the system has an invariance property: if a $j$-th coordinate of the initial vector is zero ($p_{j}^{0}=0$), then this coordinate will remain zero throughout the entire evolution of the system ($p_{j}^{t}=0$ for all $t\geq 0$). This means the corresponding component does not participate in the competition for resources. Therefore, before starting the calculations, it is necessary to filter out such indices and work only with those corresponding to non-zero initial coordinates.
$M_{0}=\{k\in\{1,\dots,n\}\mid p_{k}^{0}>0\}$
All subsequent steps are performed only for the indices $k\in M_{0}$ and their corresponding constants $c_{k}$.
2. Find the equilibrium set $M^{*}$ through iteration. This step determines the final set of indices $M^{*}\subseteq M_{0}$ that will have non-zero coordinates in the limit state. The procedure is based on the self-consistent condition from Theorem 4: a coordinate $p_{k}^{\infty}$ can be positive only if its corresponding constant $c_{k}$ exceeds a critical threshold $\Lambda(M^{*})$, which in turn depends on the set $M^{*}$ itself.
  1. 1.
    Initialization: Assume that all initially active coordinates will remain active. The candidate set $M_{cand}$ is initialized as $M_{0}$.
  2. 2.
    Iterative Loop: Repeat the following steps until the set $M_{cand}$ no longer changes:
    • 2.1)
      Compute the threshold $\Lambda(M_{cand})$: If $|M_{cand}|\leq 1$, then $\Lambda(M_{cand})=0$. Otherwise, the threshold is calculated using the formula:
      $\Lambda(M_{cand})=\frac{|M_{cand}|-1}{\sum_{k\in M_{cand}}\frac{1}{c_{k}}}$
    • 2.2)
      Update the set $M_{cand}$: Form a new version of the set, keeping only those indices $k$ for which the condition $c_{k}>\Lambda(M_{cand})$ is satisfied:
      $M_{cand}=\{k\in M_{cand}\mid c_{k}>\Lambda(M_{cand})\}$
  3. 3.
    Result: When the set $M_{cand}$ remains unchanged after an iteration, the loop terminates. This stabilized set is the desired equilibrium set $M^{*}$.
3. Calculate the coordinates of the fixed point $\bf p$$^{\infty}$. Once the set $M^{*}$ and the final value $\Lambda(M^{*})$ are known, the coordinates of the vector $\bf p$$^{\infty}$ are calculated analytically:
  • For each index $k$ belonging to the equilibrium set ($k\in M^{*}$), the value of the coordinate is calculated by the formula:
    $p_{k}^{\infty}=1-\frac{\Lambda(M^{*})}{c_{k}}$
  • For each index $j$ not belonging to the equilibrium set ($j\notin M^{*}$), the value of the coordinate is zero:
    $p_{j}^{\infty}=0$
The result of performing these steps is the vector $\bf p$$^{\infty}$, which represents the unique attractor of the system for the given parameters.
3.2 Analysis of the stability of limit states
@@
At the fixed point $p^{\infty}$, the condition $c_{i}(1-p_{i}^{\infty})=L_{c}^{\infty}$ holds for all $i$ where $p_{i}^{\infty}>0$, and $L_{c}^{\infty}$ is
@@
Since $|f^{\prime}(p_{1}^{\infty})|<1$, the fixed point $p^{\infty}$ is locally asymptotically stable.
@@
Since $c_{1},c_{2}>0$, the modulus of the derivative at both points is strictly greater than 1. This proves that the boundary fixed points $(1;0)$ and $(0,1)$ are unstable repellers. Any small perturbation away from these ”pure” states will cause the system to evolve towards the internal stable fixed point $\bf p$$^{\infty}$.
@@
Case $n\geq 3$. In the stability analysis for $n\geq 3$, we focus on initial state vectors ${\bf p}^{t=0}$ with strictly positive coordinates, i.e. points in the interior of the (n1)(n-1)-dimensional simplex. If some coordinate of the initial vector ${\bf p}^{t=0}$ is exactly zero, the trajectory remains forever on the corresponding boundary face in the absence of perturbations, since this face is an invariant set of the dynamics. However, any arbitrarily small transversal (out-of-face) perturbation makes this coordinate positive and shifts the system’s state into a higher-dimensional face of the simplex. The limiting state $p^{\infty}$ in this perturbed case will, in general, differ from the boundary equilibrium, which implies that such a boundary fixed point is unstable with respect to transversal perturbations, even if it is stable with respect to perturbations tangential to the face.
@@
A direct analytical proof of local asymptotic stability for the general case of $n\geq 3$ is infeasible. The stability of the system is determined by the eigenvalues $\lambda$ of the Jacobian matrix $J(\mathbf{p}^{\infty})$, which are the roots of the characteristic equation $\det(J(\mathbf{p}^{\infty})-\lambda I)=0$. The problem is that the elements of the matrix $J(\mathbf{p}^{\infty})$ depend on the coordinates of the fixed point $p_{i}^{\infty}$ and the value $\Lambda$. These quantities, in turn, are highly complex nonlinear functions of the initial system parameters $c_{i}$, $i=\overline{1,n}$. As a result, the coefficients of the characteristic polynomial become exceedingly cumbersome rational expressions of $c_{i}$, $i=\overline{1,n}$. The characteristic equation is a polynomial of degree $n$. According to the Abel-Ruffini theorem, for polynomials of degree $n\geq 5$, no general algebraic solution exists that expresses the roots in terms of the coefficients using a finite number of arithmetic operations and root extractions.
Given the complexity of an analytical proof, numerical methods are employed to verify stability in each specific case. To automate this process, a dedicated software tool was developed in Python (see Code Availability, stability_analysis.py) that performs a stability analysis for any given point $p^{\infty}$ and set of parameters $c_{i}$. This application supports arbitrary dimensions $n\geq 2$ and any admissible set of constants $c_{i}$, $i=\overline{1,n}$. The procedure consists of the following steps:
@@
    For a given point $p^{\infty}$, the set of indices corresponding to non-zero (active) components and the set of indices corresponding to zero components are determined. This division is key, as stability is analyzed separately for perturbations along the active directions (tangential stability) and perpendicular to them (transversal stability).
@@
    For the subspace defined by the active components, the corresponding Jacobian sub-matrix $J_{\text{active}}(p^{\infty})$ is formed. Its eigenvalues $\lambda$ are calculated, and the spectral radius $\rho(J_{\text{active}})=\max|\lambda_{i}|$ is determined. Stability against tangential perturbations is determined by the criterion $\rho(J_{\text{active}})<1$.
@@
    For each zero component $p_{i}^{\infty}=0$, the corresponding ”normal” eigenvalue $\lambda_{\text{norm},i}$ is calculated. If at least one of these values exceeds 1 ($\lambda_{\text{norm},i}>1$), the point is considered unstable to perturbations that move it out of the subspace in which it lies.
@@
    The final conclusion about the stability of the point $p^{\infty}$ is based on the results of both analyses. A point is stable only if it is stable to both tangential and transversal perturbations. The presence of transversal instability is a sufficient condition for the overall instability of the point, regardless of the result of the tangential stability analysis.
Extensive numerical experiments confirm that for a vector ${\bf p}^{t=0}$ with positive coordinates and randomly generated $c_{i}$, $i=\overline{1,n}$ satisfying the conditions of Theorems 3 and 4, the spectral radius remains below unity, thus supporting the local asymptotic stability of the computed fixed point. In cases where one or more $c_{i}$ cross the critical threshold $\Lambda(M^{*})$, the program consistently detects loss of stability through the emergence of an eigenvalue with $|\lambda_{i}|>1$.

### v2

@@ 2. $2)$
    The boundary of the simplex (the set of vectors where at least one coordinate is zero) is an invariant set. All fixed points belonging to this boundary are unstable in the full state space. Specifically, vertices act as unstable repellers, while fixed points located on the higher-dimensional faces act as topological saddles.
@@
The stability analysis of fixed points for nonlinear maps, as performed here via linearization and eigenvalue analysis of the Jacobian, is a cornerstone of dynamical systems theory [27, 28]. The result that boundary equilibria are unstable repellers while the interior equilibrium acts as an attractor is a common feature in systems modeling population dynamics and competition [29].
@@
In addition to establishing stability, the analysis of the Jacobian’s eigenvalues provides an important insight into the rate of convergence to the equilibrium. As shown, all n1n-1 eigenvalues corresponding to perturbations on the simplex are equal to $\lambda=\frac{n^{2}-2}{n^{2}-1}$. This value can be expressed as $\lambda=1-\frac{1}{n^{2}-1}$, from which it is evident that as the number of system components $n$ increases, $\lambda$ approaches 1 from below. Since the decay rate of perturbations near the fixed point is determined by the magnitude $|\lambda|^{t}$, the proximity of $\lambda$ to unity signifies that the convergence is very slow. Consequently, for systems with a large number of interacting components (a large $n$), the process of reaching a uniform distribution can be extremely prolonged. In practice, this means the system can spend long periods in transient, non-uniform states before reaching its asymptotic equilibrium.
3 Dependence of the equilibrium on the set of constants.
@@
This heterogeneity in the parameters $c_{i}$ is conceptually similar to the introduction of influential agents or “zealots” in opinion dynamics, who can prevent a full consensus [30]. Whereas models of bounded confidence achieve fragmentation through interaction thresholds [25], our model generates aggregated states analytically, based on the intrinsic “attractiveness” $c_{i}$ of each component.
@@
Next, let’s consider the case where all parameters $c_{i}$ are different. Mathematically, the equilibrium is entirely determined by this set of parameters $c_{i}$. For instance, a violation of condition (11) for some index $i$ in Theorem 3 leads to the corresponding coordinate vanishing, i.e., $p_{i}^{\infty}=0$. This behavior has been confirmed by numerous computer simulations (see Fig. 2a, 2b). Furthermore, the explicit form of the vectors of equilibrium demonstrates that a greater value of the constant $c_{i}$ strictly correlates with a greater value of the corresponding $i$-th coordinate.
In terms of the applied model, this mathematical dependence illustrates how external parameters (interpreted as assistance or favorability) govern the resource distribution [31]. Specifically, more favorable conditions associated with a certain state $\omega_{i}$ lead to an increase in the corresponding component’s share in that state. Conversely, the condition $p_{i}^{\infty}=0$ is interpreted as the complete depletion of the resource from the $i$-th state, resulting in the component’s displacement from the system.
@@
Hereafter, we will use the notation $\kappa_{i}^{t}:=c_{i}(1-p_{i}^{t})$.
@@
$V_{n}=\{{\bf e}_{1},{\bf e}_{2},\ldots,{\bf e}_{n}\},$
where ${\bf e}_{i}=(0,\ldots,1,\ldots,0)$ is the vector with a 1 in the $i$-th position and zeros elsewhere.
@@
Let the initial state of the system be ${\bf p}^{0}={\bf e}_{i}$ for some $i\in\{1,\ldots,n\}$. This means the initial components of the system are $p_{i}^{0}=1$ and $p_{j}^{0}=0$ for all $j\neq i$.
@@
3.1 Existence of a fixed point
@@
Let ${\bf p}^{0}\in\mathbb{R}_{+}^{2}$, ${\bf p}^{0}\neq{\bf e}_{i}$. Then each trajectory of the dynamic conflict system (1) converges to a fixed point
@@
Then from the invariance of the equilibrium, we have the system
@@
Next, the case of a three-dimensional space is considered, i.e., ${\bf p}\in\mathbb{R}_{+}^{3}$. Note that Theorem 2, as well as the following Theorem 3, can be generalized to the case ${\bf p}\in\mathbb{R}_{+}^{n}$, $n>3$. In this case, condition (11) becomes significantly more complex, and the description of the vector $\bf p$$^{\infty}$ breaks down into several non-trivial cases that require additional analysis.
Figure 2a: Phase portrait on the 2D invariant simplex within the 3D state space, illustrating the asymptotic trajectories of the stochastic vectors ${\bf p}^{t}$ under the action of conflict transformation (9). Case (a) with asymmetric interaction constants ${\bf c}=(0.3;0.4;0.25)$. The trajectories asymptotically converge to the interior attractor ${\bf p}^{\infty}=(0.322;0.4915;0.1865)$. The boundary fixed points act as saddles: $S_{12}(0.429;0.571;0)$, $S_{23}(0;0.615;0.385)$, and $S_{13}(0.545;0;0.455)$, with their unstable invariant manifolds defining the principal pathways of the flow toward the interior attractor. The vertices $P_{1}(1;0;0)$, $P_{2}(0;1;0)$, and $P_{3}(0;0;1)$ are unstable repelling fixed points, and all other points within the simplex represent transient states.
@@
Let ${\bf p}^{0}\in\mathbb{R}_{+}^{3}$, ${\bf p}^{0}\neq{\bf e}_{i}$, and for each $i=1,\ldots,3$ one of the conditions $\kappa^{t}_{i}>L_{c}^{t}$ or $\kappa^{t}_{i}<L_{c}^{t}$ holds for any $t$. Then every trajectory of the dynamic conflict system (1) converges to a fixed point
@@
Figure 2b: Phase portrait on the 2D invariant simplex within the 3D state space. Case (b) with ${\bf c}=(0.8;0.1;0.9)$. The condition for the existence of an interior fixed point is violated ($c_{2}<\frac{c_{1}c_{3}}{c_{1}+c_{3}}$}), inducing a transcritical bifurcation. Consequently, the boundary fixed point exchanges stability with the former interior equilibrium, becoming the interior attractor $S_{13}$ corresponding to ${\bf p}^{\infty}=(0.471;0;0.529)$, which implies the asymptotic vanishing of the second coordinate ($p_{2}^{\infty}=0$). The remaining saddles $S_{12}(0.889;0.111;0)$ and $S_{23}(0;0.1;0.9)$ generate unstable invariant manifolds that guide the entire phase flow directly into $S_{13}$. As in case (a), the vertices are unstable repelling fixed points.
@@
This follows directly from equations (9). If this condition holds for all $t$, then the sequence $\{p_{i}^{t}\}_{t=0}^{\infty}$ is monotonic. The boundedness of this sequence follows from the problem statement. Therefore, an asymptotic state $\bf p$$^{\infty}$ exists. Assume all coordinates $p_{i}^{\infty}$ are non-zero. Then the invariance of the equilibrium means that $k_{i,c}^{t}=1$ for all $i=1,2,3$. That is, we have the system
@@
It should be noted that the equilibrium is stable, as it is determined only by a fixed set of constant parameters $c_{i}$, which in turn are given at $t=0$ and do not depend on time $t$.
@@
Let $M^{*}=\{k:p_{k}^{\infty}>0\}$ be the set of indices for which the coordinates of the equilibrium vector are strictly positive. Then $\gamma(M^{*})$ is the number of elements in the set $M^{*}$. The existence and uniqueness of $M^{*}$ are determined by the iterative procedure (1). Let’s denote
@@
Let ${\bf p}\in\mathbb{R}^{n}$, $n>3$. Suppose that for each $i=1,\ldots,n$, starting from some time $t$, one of the conditions $k_{i}^{t}>L_{c}^{t}$ or $k_{i}^{t}<L_{c}^{t}$ holds. Then every trajectory of the dynamic conflict system (1) converges to a fixed point
@@
For any $i=1,\ldots,n$, one of the following equalities holds:
@@
Similarly to the proof of Theorem 3 for $n=3$, based on the monotonicity of the sequences $\{p_{i}^{t}\}$, the trajectories ${\bf p}^{t}$ converge to an equilibrium vector $\bf p$$^{\infty}$, which is a fixed point of the system.
For a fixed point $\bf p$$^{\infty}$, if $p_{i}^{\infty}>0$ (i.e., $i\in M^{*}$) and $c_{i}>0$, it follows from the dynamical equation (9) $p_{i}^{t+1}=p_{i}^{t}\cdot k_{i,c}^{t}$, where $k_{i,c}^{t}=\frac{n-1+c_{i}(1-p_{i}^{t})}{n-1+L_{c}^{t}}$, that in the stationary state $k_{i,c}^{\infty}=1$. This means that the numerator and denominator of $k_{i,c}^{\infty}$ are equal:
@@
The sum of all coordinates of the vector $\bf p$$^{\infty}$ is equal to 1. Since $p_{k}^{\infty}=0$ for $k\notin M^{*}$, then
@@
For $j\notin M^{*}$, $p_{j}^{\infty}=0$, which holds under the condition $c_{j}\leq\Lambda(M^{*})$. Therefore, the explicit form of the equilibrium depends on the set $M^{*}$, which, in turn, is determined by the entire set of constants $c_{k}$. ∎
@@
holds for all $k=1,\ldots,n$, then $\gamma(M^{*})=n$,
@@
As the mean–field vector ${\bf r}^{t}$ is not an independent variable, being defined at each step as a function of ${\bf p}^{t}$ by formula (3), its equilibrium $\bf r$$^{\infty}$ is also entirely dependent on $\bf p$$^{\infty}$. Therefore, after the coordinates $p_{i}^{\infty}$ have been determined from Theorems 24, the corresponding coordinates $r_{i}^{\infty}$ can be readily computed as:
@@
Below, we present the formal mathematical procedure to analytically compute the fixed point $\bf p$$^{\infty}$ for a given initial vector ${\bf p}^{t=0}$ and a set of control parameters $c_{i}$, $i=1,\ldots,n$. The complete programmatic implementation of this algorithm, available as find_fixed_point.py, can be found in the public repository (see Code Availability).
Input: Initial state vector ${\bf p}^{0}\in\mathbb{R}^{n}_{+}$, control parameters $c_{1},\ldots,c_{n}>0$.
Output: The equilibrium vector ${\bf p}^{\infty}\in\mathbb{R}^{n}_{+}$.
Step 1.
Initialization (Filtering trivial components): Determine the initial set of active components that participate in the competition:
$M_{0}=\{k\in\{1,\ldots,n\}\mid p_{k}^{0}>0\}.$
Set the initial candidate set $M_{cand}=M_{0}$.
Step 2.
Iterative thresholding (Finding the equilibrium set $M^{*}$): Update the candidate set $M_{cand}$ iteratively:
  1. (2.a)
    Calculate the critical threshold $\Lambda(M_{cand})$ for the current set: Λ(M_cand) = {0, if —Mcand— ≤1,—Mcand— - 1k ∈Mcand1ck, otherwise.
  2. (2.b)
    Identify the subset of indices that satisfy the strict survival condition: M_new = {k ∈M_cand ∣c_k ¿ Λ(M_cand)}.
  3. (2.c)
    Check for stabilization: if $M_{new}=M_{cand}$, the iteration terminates. Set the final equilibrium set $M^{*}=M_{cand}$ and proceed to Step 3. Otherwise, update $M_{cand}=M_{new}$ and repeat Step 2.
Step 3.
Computation of the final state: Calculate the exact coordinates of the fixed point $\bf p$$^{\infty}$ for all $i=1,\ldots,n$:
$p_{i}^{\infty}=\begin{cases}1-\frac{\Lambda(M^{*})}{c_{i}},&\text{if }i\in M^{*},\\ 0,&\text{if }i\notin M^{*}.\end{cases}$
3.2 Analysis of the stability of equilibria
@@
At the fixed point $\bf p$$^{\infty}$, the condition $c_{i}(1-p_{i}^{\infty})=L_{c}^{\infty}$ holds for all $i$ where $p_{i}^{\infty}>0$, and $L_{c}^{\infty}$ is
@@
Since $|f^{\prime}(p_{1}^{\infty})|<1$, the fixed point $\bf p$$^{\infty}$ is locally asymptotically stable.
@@
Since $c_{1},c_{2}>0$, the modulus of the derivative at both points is strictly greater than 1. This proves that the boundary fixed points $(1;0)$ and $(0,1)$ are unstable repellers. Any small perturbation away from these “pure” states will cause the system to evolve towards the internal stable fixed point $\bf p$$^{\infty}$.
@@
Case $n\geq 3$. In the stability analysis for $n\geq 3$, we focus on initial state vectors ${\bf p}^{t=0}$ with strictly positive coordinates, i.e. points in the interior of the (n1)(n-1)-dimensional simplex. If some coordinate of the initial vector ${\bf p}^{t=0}$ is exactly zero, the trajectory remains forever on the corresponding boundary face in the absence of perturbations, since this face is an invariant set of the dynamics. However, any arbitrarily small transversal (out-of-face) perturbation makes this coordinate positive and shifts the system’s state into a higher-dimensional face of the simplex. The equilibrium $\bf p$$^{\infty}$ in this perturbed case will, in general, differ from the boundary equilibrium, which implies that such a boundary fixed point is unstable with respect to transversal perturbations, even if it is stable with respect to perturbations tangential to the face.
@@
A direct analytical proof of local asymptotic stability for the general case of $n\geq 3$ is infeasible. The stability of the system is determined by the eigenvalues $\lambda$ of the Jacobian matrix $J(\mathbf{p}^{\infty})$, which are the roots of the characteristic equation $\det(J(\mathbf{p}^{\infty})-\lambda I)=0$. The problem is that the elements of the matrix $J(\mathbf{p}^{\infty})$ depend on the coordinates of the fixed point $p_{i}^{\infty}$ and the value $\Lambda$. These quantities, in turn, are highly complex nonlinear functions of the initial system parameters $c_{i}$, $i=1,\ldots,n$. As a result, the coefficients of the characteristic polynomial become exceedingly cumbersome rational expressions of $c_{i}$, $i=1,\ldots,n$. The characteristic equation is a polynomial of degree $n$. According to the Abel-Ruffini theorem, for polynomials of degree $n\geq 5$, no general algebraic solution exists that expresses the roots in terms of the coefficients using a finite number of arithmetic operations and root extractions.
Given the complexity of an analytical proof, numerical methods are employed to verify stability in each specific case. To automate this process, a dedicated software tool was developed in Python (see Code Availability, stability_analysis.py) that performs a stability analysis for any given point $\bf p$$^{\infty}$ and set of parameters $c_{i}$. This application supports arbitrary dimensions $n\geq 2$ and any admissible set of constants $c_{i}$, $i=1,\ldots,n$. The procedure consists of the following steps:
@@
    For a given point $\bf p$$^{\infty}$, the set of indices corresponding to non-zero (active) components and the set of indices corresponding to zero components are determined. This division is key, as stability is analyzed separately for perturbations along the active directions (tangential stability) and perpendicular to them (transversal stability).
@@
    For the subspace defined by the active components, the corresponding Jacobian sub-matrix $J_{\text{active}}({\bf p}^{\infty})$ is formed. Its eigenvalues $\lambda$ are calculated, and the spectral radius $\rho(J_{\text{active}})=\max|\lambda_{i}|$ is determined. Stability against tangential perturbations is determined by the criterion $\rho(J_{\text{active}})<1$.
@@
    For each zero component $p_{i}^{\infty}=0$, the corresponding “normal” eigenvalue $\lambda_{\text{norm},i}$ is calculated. If at least one of these values exceeds 1 ($\lambda_{\text{norm},i}>1$), the point is considered unstable to perturbations that move it out of the subspace in which it lies.
@@
    The final conclusion about the stability of the point $\bf p$$^{\infty}$ is based on the results of both analyses. A point is stable only if it is stable to both tangential and transversal perturbations. The presence of transversal instability is a sufficient condition for the overall instability of the point, regardless of the result of the tangential stability analysis.
Extensive numerical experiments confirm that for a vector ${\bf p}^{t=0}$ with positive coordinates and randomly generated $c_{i}$, $i=1,\ldots,n$ satisfying the conditions of Theorems 3 and 4, the spectral radius remains below unity, thus supporting the local asymptotic stability of the computed fixed point. In cases where one or more $c_{i}$ cross the critical threshold $\Lambda(M^{*})$, the program consistently detects loss of stability through the emergence of an eigenvalue with $|\lambda_{i}|>1$.